109/36=35/12+1/4a
a=?

Answer by tommyt3rd(5050)

109/36=35/12+1/4a
1/4 a = 109/36-35/12
a = 4(109/36-35/12) =
4(109/36-105/36) =
4(4/36)=
4/9